write the equation of a line with the given properties, in slope -intercept form
slope= 5
-
6 x , through (7,3)

Answer by algebrapro18(249)

There are 2 ways to do this problem.

Method 1: Point-slope formula.

To be able to use the point-slope formula we need both the slope of the line and a point on the line. Since we have both we can use it and we just plug and chug.

y-y1=m(x-x1) where m is the slope, and x1,y1 is the point given.

y-3=5/6(x-7)
y-3=5/6x-35/6
y=5/6x-35/6+3
y=5/6x-17/6

Method 2: Slope-intercept formula.

The other method we can use to find the equation of the line is to plug in the given information to the slope-intercept formula and then solve for b.

y=mx+b
3 = 5/6(7) + b
3 = 35/6 + b
-17/6 = b

y = 5/6x-17/6